# Clear Nights Ahead: Towards Multi-Weather Nighttime Image Restoration

### v1

@@ 3.1 Dataset Curation
On the other hand, nighttime images often suffer from light flares, which are further exacerbated and intertwined with adverse weather effects. Therefore, we also introduce more flare effects into the clear images to faithfully simulate real-world nighttime conditions:
$\mathbf{X}_{\textit{flare}}=\alpha\cdot\mathbf{X}+\beta\cdot(\mathbf{L}\ast\mathrm{K}_{\textit{APSF}}),\vspace{-0.5em}$ (2)
where $\mathbf{X}_{\textit{flare}}$ is the output flare-degraded image. $\alpha$ is a predefined hyper-parameter for clean image preservation, and $\beta$ is a dynamic hyper-parameter for flare content blending, determined by the pixel ratio of the light source region. The light source map $\mathbf{L}$ is extracted from the clean image with thresholding and alpha matting refinement [28]. $\mathrm{K}_{\textit{APSF}}$ is a 2D kernel derived from the atmospheric point spread function as introduced in [20], and $\ast$ is the convolution operator.
3.2 Dataset Analysis
Impact of Illumination-Aware Synthesis. To demonstrate the effectiveness of the proposed illumination-aware degradation generation, we analyze four different synthetic image variants of nighttime scenes (showcased in Figure 3). Ten volunteers are first recruited to rank image quality along three dimensions (i.e., authenticity, lighting complexity and task relevance). As illustrated in Figure 4, the combination of illumination-aware weather and flare achieves the best human preference. We evaluate whether training with our illumination-aware synthesized data improves image restoration performance in real-world scenarios. Figure 4 demonstrates that four representative models [68, 8, 65] trained with our synthesized data exhibit clear and consistent improvements. The images synthesized using our illumination-aware generation method not only exhibit greater visual realism, but also encourage models to disentangle these complex degradations, thereby enhancing restoration capabilities in challenging nighttime scenarios.
Figure 5: Distribution of the introduced AllWeatherNight dataset.
Dataset Statistics. Overall, we synthesize 8,000 nighttime images for model training, covering both multi- and single-degradation scenarios with varying scales, directions, patterns and intensities. As summarized in Figure 5, the test dataset comprises two subsets (i.e., synthetic and real-world subsets), each containing 1,000 images. The synthetic subset evaluates models across seven scenes, comprising both compound and single degradations. The real-world subset is categorized into four types of adverse weather conditions, enabling model evaluation in natural scenarios.
4 ClearNight
Figure 6: Overview of our ClearNight framework. ClearNight primarily comprises Retinex-based dual prior guidance as well as weather-aware dynamic specificity and commonality branches. The extracted Retinex-based dual priors explicitly guide the network to focus on uneven illumination regions and intrinsic texture contents respectively. The weather-aware dynamic specificity branch adaptively accommodates specific weather conditions and collaborates with the commonality branch to effectively address multiple weather degradations.
Multi-weather nighttime image restoration aims to recover clear nighttime images degraded by flare effects and multiple adverse weather conditions (e.g., haze, rain, snow and the compound degradations). This paper introduces ClearNight, a unified nighttime image restoration framework that can remove multiple weather degradations in one go. As depicted in Figure 6, ClearNight integrates Retinex-based dual prior guidance and weather-aware dynamic specificity-commonality collaboration. We explicitly decouple the uneven lighting and texture information, and extract Retinex-based dual priors to guide the network in recovering clear images with natural lighting and rich background details. Building on the Swin-Unet-like architecture [43, 2], we devise a weather-aware dynamic specificity branch and establish its collaboration with a commonality branch to exploit the unique and shared characters of diverse weather degradations, enhancing the capability to restore images affected by multiple adverse weather conditions.
4.1 Retinex-Based Dual Prior Guidance
Decoupling illumination and texture information is critical for nighttime image restoration, as it enables effective handling of non-uniform lighting and adverse weather effects. Guided by this insight, we extract multi-scale Retinex-based dual priors to explicitly guide the restoration processes, as shown in Figure 6. Specifically, based on Retinex theory [10, 27], we decompose input degraded image $\mathbf{X}_{\textit{d}}$:
$\mathbf{X}_{\textit{d}}=\mathbf{R}_{\textit{d}}\cdot\mathbf{I}_{\textit{d}},\vspace{-0.5em}$ (3)
where $\mathbf{R}_{\textit{d}}$ and $\mathbf{I}_{\textit{d}}$ are the reflectance and illumination components decomposed from the input degraded image. The illumination component highlights the non-uniform lighting conditions, while the reflectance component represents intrinsic texture information of the image content.
The two decomposed components $\mathbf{p}^{\textit{rtx}}\in\{\mathbf{I}_{\textit{d}},\mathbf{R}_{\textit{d}}\}$ are then fed into independent multi-scale prior extraction units (MPE) to extract Retinex-based dual priors at three scales:
$\mathbf{p}^{1},\mathbf{p}^{2},\mathbf{p}^{3}=\mathrm{MPE}(\mathbf{p}^{\textit{rtx}})\ \ \quad\textrm{with}\ \ \quad\mathbf{p}^{n}\in\{\mathbf{i}^{n},\mathbf{r}^{n}\},\vspace{-0.5em}$ (4)
where $\mathbf{p}^{n}$ denotes the extracted illumination or reflectance prior at the $n$-th scale, with $n\in\{1,2,3\}$. The dual priors are extracted using MPE with identical structures but without weight sharing. The detailed architecture of MPE is depicted in the top-middle of Figure 6. A dilated convolution is employed to project $\mathbf{p}^{\textit{rtx}}$ into three scales, then the three-scale features are interactively fused to produce the final multi-scale illumination/reflectance priors $\mathbf{p}^{n}$. These multi-scale priors are subsequently used to guide the network, ensuring the awareness of illumination regions and intrinsic patterns across different stages of the network.
More specifically, the extracted illumination priors $\mathbf{i}^{n}$ are successively injected into the Transformer blocks (TFBs) of the first three stages, which guide the network to decouple illumination from textures and pay more attention to the uneven lighting regions in nighttime images, thereby facilitating the handling of lighting-influenced weather degradations. As the reflectance priors $\mathbf{r}^{n}$ contain rich intrinsic textures that not only capture background details but also reveal degradation types, we incorporate them into each weather-aware dynamic selection block (WDS) in the specificity branch to enhance weather type discrimination and improve multi-weather restoration performance.
Figure 7: Qualitative results on AllWeatherNight synthetic testing subset. Please zoom in for a better view.
4.2 Weather-Aware Dynamic Specificity and Commonality Collaboration
Dynamic Specificity and Commonality Synergy. As different weather degradations (e.g., snow and rain streaks) share certain commonalities while exhibiting unique characteristics, we introduce a synergistic design of dynamic specificity and commonality branches to effectively model complex multi-weather degradations. The commonality branch is built upon [43] consisting of sequential Transformer blocks with skip connections, while the specificity branch incorporates multiple weather-aware dynamic selection blocks (WDS) as residuals to each stage of TFBs.
As shown in the top-right of Figure 6, each WDS consists of an encoder-decoder structure coupled with a dynamic selection module. The WDS block merges the interacted features of the two branches and the reflectance priors as input. Inspired by [66, 16], we adopt the encoder-decoder structure to construct a compact feature space that benefits WDS in learning discriminative weather-specific representation. The dynamic selection module contains multiple candidate units $\mathrm{U}_{k}(\cdot)$ and learns to dynamically construct distinct sub-networks by sparsely selecting candidates tailored to different inputs, where $k$ is the unit index.
Besides the candidate units, the dynamic selection module includes a gating layer $\mathrm{W}(\cdot)$ that computes the importance weights assigned to each candidate unit, and a router $\mathrm{R}(\cdot)$ that estimates the probability of each candidate unit being selected. Given the input feature $\bar{\mathbf{f}}^{a}$, the output of the $a$-th dynamic selection module can be calculated by
$\!\mathbf{\bar{M}}^{a}\!\!=\!\!\sum_{k\in\mathcal{T}}\mathrm{W}(\bar{\mathbf{f}}^{a})_{k}\cdot\mathrm{U}_{k}(\bar{\mathbf{f}}^{a})\quad\textrm{with}\quad\mathcal{T}\!\!=\!\!\mathrm{TopK}(\mathrm{R}(\bar{\mathbf{f}}^{a})),\!\vspace{-0.5em}$ (5)
where $\mathcal{T}$ denotes the set of selected top-$K$ unit indices, and $\mathrm{W}(\bar{\mathbf{f}}^{a})_{k}$ is the importance weight for the $k$-th unit.
Dynamic Weather Degradation Modeling. The dynamic selection module adaptively selects candidate units based solely on the visual content of each input sample, which fails to fully capture the correlations and distinctive characteristics of different weather types, thereby limiting its effectiveness in handling challenging multi-weather scenarios. To address this, we introduce a weather instructor to associate distinct weather types with designated candidate units, enabling more structured modeling and improving generalization across diverse weather combinations.
Specifically, we reorganize the dynamic selection module into distinct parts: the candidate units, router and gating layer serve as dynamic allocators, while the shared linear layers preceding them functions as the knowledge-sharing unit. We further introduce a classifier after the knowledge-sharing unit as weather instructor to aggregate features corresponding to the same weather condition by learning a set of prototypes (one for each single weather degradation):
$y^{a}=\mathrm{WI}(\bar{\mathbf{f}}^{a})\ \ \quad\textrm{and}\ \ \quad\bar{\mathbf{f}}^{a}=\mathrm{KSU}(\mathbf{f}^{a}_{\textit{enc}}),\vspace{-0.5em}$ (6)
where $\mathrm{KSU}(\cdot)$ and $\mathrm{WI}(\cdot)$ denote the knowledge-sharing unit and weather instructor, respectively. $\mathbf{f}^{a}_{\textit{enc}}$ indicates the output of the encoder, and $y^{a}$ is the weather type predicted by the $a$-th dynamic selection module.
The weather instructor performs multi-label classification on each input, supervised by a binary cross-entropy loss $\mathcal{L}_{\textit{bce}}$ with respect to the corresponding weather tags. During training, these prototypes pull the features from the corresponding weather towards themselves. Under the guidance of the weather instructor, features $\bar{\mathbf{f}}^{a}$ of the images affected by the same weather degradations tend to share similar representations. Consequently, the dynamic allocator is implicitly guided to be weather-aware and consistently selects similar candidate units for the inputs affected by the same weather degradations, thereby improving the diversity and specificity of candidate units and restoration effectiveness.
Figure 8: Qualitative results on AllWeatherNight real-world testing subset. Please zoom in for a better view.
Table 2: Quantitative results on AllWeatherNight synthetic testing subset. The best and second-best results are highlighted in red and blue.
Multi-Degradation Single-Degradation
Method Rain Scene Snow Scene Haze Rain Streak Raindrop Snow Flare
@@ 4.2 Weather-Aware Dynamic Specificity and Commonality Collaboration
TKL [6] 29.0919 / 0.8769 26.2657 / 0.8456 31.6136 / 0.9401 30.4227 / 0.8844 32.5573 / 0.9561 31.0471 / 0.9247 36.7239 / 0.9741
FSDGN [57] 29.9850 / 0.8730 28.0147 / 0.8474 34.1807 / 0.9378 31.3309 / 0.8851 33.9535 / 0.9638 31.9022 / 0.9288 38.7223 / 0.9798
WeatherDiff [69] 29.7631 / 0.8936 27.1109 / 0.8537 30.5020 / 0.9256 33.3630 / 0.9352 35.9385 / 0.9726 33.6573 / 0.9488 40.2057 / 0.9826
WGWS [68] 30.5811 / 0.8961 27.9023 / 0.8658 31.6132 / 0.9181 32.8077 / 0.9221 34.3616 / 0.9554 32.1325 / 0.9196 32.1085 / 0.9017
RLP [60] 21.2180 / 0.6641 19.3059 / 0.6124 18.6348 / 0.6379 31.1586 / 0.8867 32.4204 / 0.9322 31.5645 / 0.9101 34.7658 / 0.9451
RAMiT [8] 30.4565 / 0.9106 29.1169 / 0.8889 36.4414 / 0.9738 31.9433 / 0.9204 33.8452 / 0.9632 32.8934 / 0.9491 43.0080 / 0.9934
DiffUIR [65] 27.6676 / 0.8040 25.8151 / 0.7892 28.4763 / 0.8341 30.2547 / 0.8624 29.6778 / 0.8740 26.9833 / 0.8055 31.2789 / 0.8816
DEA-Net [7] 31.4244 / 0.9202 29.2500 / 0.8956 35.8174 / 0.9612 32.7631 / 0.9285 34.8406 / 0.9704 33.5811 / 0.9493 38.6533 / 0.9807
@@
ClearNight (ours) 32.5937 / 0.9223 30.6464 / 0.9041 36.4655 / 0.9621 33.6238 / 0.9331 35.4282 / 0.9723 33.9747 / 0.9539 38.7707 / 0.9838
4.3 Network Optimization
@@
Multiple losses are utilized to jointly optimize ClearNight, where L1 loss $\mathcal{L}_{\textit{1}}$ and perceptual loss $\mathcal{L}_{\textit{per}}$ [22] are used to ensure that restored results resemble ground-truths closely. To enhance the structure and details of outputs, depth loss $\mathcal{L}_{\textit{depth}}$ is exploited via minimizing the differences between ground-truths and predictions of a pre-trained depth estimation model [32]. In addition, we use a load balancing loss $\mathcal{L}_{\textit{lb}}$ [41] to balance the utilization of the candidate units. We jointly optimize the network using the total loss function:
$\mathcal{L}_{\textit{total}}=\mathcal{L}_{\textit{1}}+\lambda_{\textit{per}}\mathcal{L}_{\textit{per}}+\lambda_{\textit{bce}}\mathcal{L}_{\textit{bce}}+\lambda_{\textit{lb}}\mathcal{L}_{\textit{lb}}+\lambda_{\textit{depth}}\mathcal{L}_{\textit{depth}},\vspace{-0.5em}$ (7)
where $\lambda_{\textit{per}}$, $\lambda_{\textit{bce}}$, $\lambda_{\textit{lb}}$ and $\lambda_{\textit{depth}}$ are the loss weights.
5 Experiments
5.1 Implementation Details
@@
The synthetic image size is 640 $\times$ 360. During training, the input image is randomly cropped to 256 $\times$ 256. We adopt DehazeFormer [43] as our baseline for its advanced normalization and spatial aggregation, allowing effective natural lighting restoration. We use Adam [25] as optimizer and the initial learning rate is set to $2\times 10^{-4}$ for 100 epochs. The learning rate is adjusted using the cosine annealing scheme [18]. The loss weight $\lambda_{\textit{per}}$, $\lambda_{\textit{bce}}$, $\lambda_{\textit{lb}}$ and $\lambda_{\textit{depth}}$ are set to 0.1, 0.001, 0.01 and 0.02, respectively. $\alpha$ is set to 0.995.
@@
Method Haze Rain Streak Raindrop Snow
TKL [6] 4.1872 3.7765 3.9238 3.2680
FSDGN [57] 4.2780 4.4694 4.9149 4.1528
WeatherDiff [69] 4.1964 3.7842 3.9254 3.3451
WGWS [68] 4.1879 3.7732 3.9635 3.2769
RLP [60] 4.9699 4.1882 5.6240 4.7669
RAMiT [8] 4.1655 3.9298 4.0808 3.2497
DiffUIR [65] 4.1063 3.8728 4.0471 3.3547
DEA-Net [7] 4.1665 3.9826 4.0889 3.4201
@@
ClearNight (ours) 4.1623 3.7653 3.8882 3.2191
5.2 Comparison with the State-of-the-Arts
Results on Synthetic Data. We compare ClearNight with state-of-the-art adverse weather image restoration methods on AllWeatherNight, including multi-weather image restoration method [6, 69, 68, 8, 65] and task-specific approaches [60, 7, 57]. As shown in Figure 7, TKL [6], FSDGN [57] and DiffUIR [65] produce overly smooth results such as the clock and railing regions, while other approaches often lose structural details or leave residual artifacts. In contrast, ClearNight preserves rich background details and restores natural lighting, illustrating robust performance across diverse and complex nighttime scenarios.
Table 2 reports quantitative results on multi-degradation and single-degradation scenes, evaluated using PSNR [15] and SSIM [67] for synthetic samples. Among these methods, FSDGN [57], RLP [60] and DEA-Net [7] are tailored for daytime/nighttime task-specific image restoration, yet they exhibit limited performance in other scenarios due to their inability to extract diverse weather features. In contrast, our method focuses on robustly addressing weather effects in nighttime scenes. As the training data lack dedicated flare samples, its performance in flare removal remains moderate. Nevertheless, ClearNight achieves the best results on the multi-degradation subset and delivers competitive performance on single-degradation samples.
Results on Real-World Data. To further validate ClearNight’s performance, we conduct qualitative and quantitative experiments on real-world samples. As shown in Figure 8, ClearNight effectively removes most weather effects and mitigates flare impacts, even when degradation blends with background content. Our method can consistently restore more natural results compared to state-of-the-art methods. Furthermore, we use NIQE [37] to assess the restored images, as shown in Table 3. The experimental results demonstrate that ClearNight can predict highly realistic images under various adverse weather conditions.
Table 4: Ablation study of key component. IP and RP denote illumination prior and reflectance prior, respectively. DA and WI indicate dynamic allocator and weather indicator in WDS.
@@ 5.2 Comparison with the State-of-the-Arts
case 1 28.7976 / 0.8825
case 2 32.1304 / 0.9176
case 3 31.7075 / 0.9113
case 4 32.2393 / 0.9179
case 5 32.4430 / 0.9214
case 6 32.2528 / 0.9184
ClearNight 32.5937 / 0.9223
Table 5: Ablation on Retinex-based dual prior extraction.
(a) impact of illumination prior
@@
28.7976 / 0.8825
31.6843 / 0.9109
31.6714 / 0.9114
31.6429 / 0.9110
31.7217 / 0.9115
32.1304 / 0.9176
(b) impact of reflection prior
[…]
31.7075 / 0.9113
31.9680 / 0.9198
31.9374 / 0.9201
31.8438 / 0.9172
32.0335 / 0.9203
32.0594 / 0.9205
5.3 Ablation Studies
@@
Impact of Model Components. We evaluate the effectiveness of the illumination prior, reflectance prior, dynamic allocator and weather indicator on the Rain Scene testing subset. As shown in Table 4, dual priors significantly enhance the restoration performance over the baseline, while the dynamic allocator and the weather indicator further optimize candidate unit selection to handle multi-weather scenes. Quantitative results demonstrate that the integration of all components achieves the best performance.
Influence of Retinex-Based Dual Prior Guidance. To evaluate the impact of multi-scale prior, we conduct ablation studies on illumination and reflectance priors at various scales, using case 1 and case 3 as baselines, respectively. As shown in Table 5, incorporating multi-scale dual priors into the network enables it to better focus on lighting and texture features, significantly enhancing restoration performance.
Effectiveness of WDS. We investigate the number of candidate units ($B$) and top-$K$ selection parameter ($K$). As shown in Figure 9, increasing the number of units enhances selective capacity, with an optimal $K$ being critical for superior performance. Notably, performance improves with larger $K$ up to a threshold, beyond which irrelevant information causes degradation.
@@
Figure 10 illustrates the relationships between various degradations and selected units, demonstrating the effectiveness of WDS in capturing the specificity of different degradations. The WDS differentiates weather features, selecting similar unit combinations for the same weather and different ones for distinct weather, enabling the network to efficiently eliminate diverse weather distortions. We visualize the feature $\bar{\mathbf{f}}^{a}$ of different weather types in Figure 11. Despite the visual similarity between raindrop and snow, ClearNight still differentiates them. Notably, the features of H+S lie between haze and snow features, indicating that our model learns the correlations among multiple weather effects in complex nighttime scenarios.
@@
Figure 10: Correlation of selected units and diverse degradations. The WDS associates various distortions with candidate units.
(a) four different single-degradations
(b) single- and multi-degradations
@@
Figure 11: Feature distribution visualization of different weather using t-SNE. H, RS, RD and S indicate haze, rain streak, raindrop and snow. The number after weather types is the index of WDS.
6 Conclusion

### v2

@@ Introduction
[…]
TKL 29.0919 / 0.8769 26.2657 / 0.8456 31.6136 / 0.9401 30.4227 / 0.8844 32.5573 / 0.9561 31.0471 / 0.9247 36.7239 / 0.9741
FSDGN 29.9850 / 0.8730 28.0147 / 0.8474 34.1807 / 0.9378 31.3309 / 0.8851 33.9535 / 0.9638 31.9022 / 0.9288 38.7223 / 0.9798
WeatherDiff 29.7631 / 0.8936 27.1109 / 0.8537 30.5020 / 0.9256 33.3630 / 0.9352 35.9385 / 0.9726 33.6573 / 0.9488 40.2057 / 0.9826
WGWS 30.5811 / 0.8961 27.9023 / 0.8658 31.6132 / 0.9181 32.8077 / 0.9221 34.3616 / 0.9554 32.1325 / 0.9196 32.1085 / 0.9017
RLP 21.2180 / 0.6641 19.3059 / 0.6124 18.6348 / 0.6379 31.1586 / 0.8867 32.4204 / 0.9322 31.5645 / 0.9101 34.7658 / 0.9451
RAMiT 30.4565 / 0.9106 29.1169 / 0.8889 36.4414 / 0.9738 31.9433 / 0.9204 33.8452 / 0.9632 32.8934 / 0.9491 43.0080 / 0.9934
DiffUIR 27.6676 / 0.8040 25.8151 / 0.7892 28.4763 / 0.8341 30.2547 / 0.8624 29.6778 / 0.8740 26.9833 / 0.8055 31.2789 / 0.8816
DEA-Net 31.4244 / 0.9202 29.2500 / 0.8956 35.8174 / 0.9612 32.7631 / 0.9285 34.8406 / 0.9704 33.5811 / 0.9493 38.6533 / 0.9807
AWRaCLe 31.5392 / 0.9210 29.4270 / 0.8738 36.4315 / 0.9599 33.1078 / 0.9317 35.2985 / 0.9686 33.6716 / 0.9532 40.1014 / 0.9836
@@
ClearNight 32.5937 / 0.9223 30.6464 / 0.9041 36.4655 / 0.9621 33.6238 / 0.9331 35.4282 / 0.9723 33.9747 / 0.9539 38.7707 / 0.9838
Table 2: Quantitative results on AllWeatherNight synthetic testing subset. The best and second-best results are highlighted.
Network Optimization
@@
Multiple losses are utilized to jointly optimize ClearNight, where L1 loss $\mathcal{L}_{\mathrm{1}}$ and perceptual loss $\mathcal{L}_{\mathrm{p}}$ (per) are used to ensure that restored results resemble ground-truths closely. To enhance the structure and details of outputs, depth loss $\mathcal{L}_{\mathrm{d}}$ is exploited via minimizing the differences between ground-truths and predictions of a pre-trained depth estimation model (depth). In addition, we use a load balancing loss $\mathcal{L}_{\mathrm{lb}}$ (losslb) to balance the utilization of the candidate units. We jointly optimize the network using the total loss function:
$\mathcal{L}_{\mathit{total}}=\mathcal{L}_{\mathrm{1}}+\lambda_{\mathrm{p}}\mathcal{L}_{\mathrm{p}}+\lambda_{\mathrm{bce}}\mathcal{L}_{\mathrm{bce}}+\lambda_{\mathrm{lb}}\mathcal{L}_{\mathrm{lb}}+\lambda_{\mathrm{d}}\mathcal{L}_{\mathrm{d}},$ (7)
where $\lambda_{\mathrm{p}}$, $\lambda_{\mathrm{bce}}$, $\lambda_{\mathrm{lb}}$ and $\lambda_{\mathrm{d}}$ are the loss weights.
Experiments
Implementation Details
@@
The synthetic image size is 640 $\times$ 360. During training, the input image is randomly cropped to 256 $\times$ 256. We adopt DehazeFormer (swin) as our baseline for its advanced spatial aggregation, allowing effective natural lighting restoration. We use Adam as optimizer and the initial learning rate is set to $2\times 10^{-4}$ for 100 epochs. The learning rate is adjusted using the cosine annealing scheme. The loss weight $\lambda_{\mathrm{p}}$, $\lambda_{\mathrm{bce}}$, $\lambda_{\mathrm{lb}}$ and $\lambda_{\mathrm{d}}$ are empirically set to 0.1, 0.001, 0.01 and 0.02, respectively. $\alpha$ is set to 0.995.
Dataset Analysis
To demonstrate the effectiveness of illumination-aware degradation generation, we analyze four different synthetic variants of nighttime images (showcased in Fig. 3). Ten volunteers are recruited to rank image quality on authenticity, lighting complexity and task relevance. As shown in the left of Fig. 6, we can observe that the combination of illumination-aware weather and flare achieves the best human preference. Furthermore, the right part of Fig. 6 demonstrates that training with our illumination-aware data improves performance in real-world scenes, with four representation models (LWGWSCVPR2023; RAMITCVPR2024; SCVPR2024) exhibiting consistent improvements.
To further validate the proposed dataset, we compare the performance of models trained on our AllWeatherNight dataset with those trained on a composite dataset constructed from existing nighttime datasets, as shown in Tab. 1. As shown in Fig. 7, models trained on AllWeatherNight exhibit superior performance on real-world samples, which is attributable to our illumination-aware synthetic images simulating degradations more realistically.
Comparison with the State-of-the-Art
Results on Synthetic Data. We compare ClearNight with state-of-the-art adverse weather image restoration methods on AllWeatherNight, including TKL (TKLCVPR2022), FSDGN (FSDGNECCV2022), RLP (NICCV2023), WeatherDiff (WeatherDiff2023), WGWS (LWGWSCVPR2023), RAMiT (RAMITCVPR2024), DiffUIR (SCVPR2024), DEA-Net (TIP2024DEA) and AWRaCLe (AAAAI2025). As shown in Fig. 8a, TKL (TKLCVPR2022), FSDGN (FSDGNECCV2022) and DiffUIR (SCVPR2024) produce overly smooth results such as the clock and railing regions, while other approaches often lose structural details or leave residual artifacts. In contrast, ClearNight preserves rich background details and restores natural lighting, illustrating robust performance across diverse and complex nighttime scenarios.
Tab. 2 reports quantitative results on multi-degradation and single-degradation scenes, evaluated using PSNR (PSNR) and SSIM (SSIM) for synthetic samples. Among these methods, FSDGN, RLP and DEA-Net are tailored for daytime/nighttime task-specific restoration, but struggle in other task scenes due to limited weather feature extraction. In contrast, our method targets robust nighttime weather effects removal. As the training data lack dedicated flare samples, its performance in flare removal remains moderate. Nevertheless, ClearNight achieves the best results on the multi-degradation subset and delivers competitive performance on single-degradation samples.
Results on Real-World Data. As shown in Fig. 8b, ClearNight effectively removes most weather effects on real images and mitigates flares, producing more natural results compared to state-of-the-art methods. Furthermore, we use NIQE (NIQE) to assess the restored images in Tab. 3. The experimental results demonstrate that ClearNight can predict highly realistic images under various adverse weather conditions.
Comparison with Cascade Solutions. As demonstrated in Fig. 9, we compare ClearNight against two pre-trained task-specific models (NAAAI2025; NICCV2023) on real-world nighttime rain streak images from AllWeatherNight. ClearNight not only achieves superior results visually and quantitatively, but requires significantly less inference time.
@@
TKL 4.1872 3.7765 3.9238 3.2680
FSDGN 4.2780 4.4694 4.9149 4.1528
WeatherDiff 4.1964 3.7842 3.9254 3.3451
WGWS 4.1879 3.7732 3.9635 3.2769
RLP 4.9699 4.1882 5.6240 4.7669
RAMiT 4.1655 3.9298 4.0808 3.2497
DiffUIR 4.1063 3.8728 4.0471 3.3547
DEA-Net 4.1665 3.9826 4.0889 3.4201
AWRaCLe 4.1681 3.9516 4.0936 3.3398
@@
ClearNight 4.1623 3.7653 3.8882 3.2191
@@
1 28.7976 / 0.8825
2 32.1304 / 0.9176
3 31.7075 / 0.9113
4 32.2393 / 0.9179
5 32.4430 / 0.9214
6 32.2528 / 0.9184
7 32.5937 / 0.9223
Table 4: Ablation study of key component. IP and RP denote illumination and reflectance priors. DA and WI indicate dynamic allocator and weather instructor in WDS.
Ablation Studies
Figure 9: ClearNight outperforms a cascade of two latest task-specific nighttime restoration methods. Quantitative results are averaged over all the real-world rain streak images.
@@
Ablation on Model Components. We evaluate the effectiveness of the illumination prior, reflectance prior, dynamic allocator and weather instructor on the Rain Scene testing subset. As shown in Tab. 4, dual priors significantly enhance the restoration performance over the baseline, while the dynamic allocator, guided by the weather instructor, optimizes candidate unit selection to handle multi-weather scenes. Quantitative results demonstrate that the integration of all components achieves the best performance.
@@
Effectiveness of WDS. Fig. 10 illustrates the relationships between various degradations and selected units, demonstrating the effectiveness of WDS in capturing the specificity of different degradations. The WDS differentiates weather features, selecting similar unit combinations for degraded scenes with the same weather, enabling the network to efficiently eliminate diverse distortions. We visualize the feature $\bar{\mathit{f}}^{a}$ of different weather types in Fig. 11. Despite the visual similarity between rain and snow, ClearNight still differentiates them. Notably, the features of $\mathrm{H+S}$” lie between haze and snow features, indicating that our model learns the correlations among multiple weather effects in complex nighttime scenarios.
@@
Figure 10: Correlation of selected units and diverse degradations. WDS associates various distortions with candidate units.
@@
Figure 11: T-SNE visualization of feature distributions across distinct weather. H, RS, RD and S indicate haze, rain streak, raindrop and snow. The number is the index of WDS.
Conclusion